# FuzzCoh: Robust Canonical Coherence-Based Fuzzy Clustering of Multivariate Time SeriesLABEL:lastpage

### v1

@@ 2.2.2 The cluster validity index for the automatic selection of $C$ and $m$
A higher FSIΩ value indicates that the fuzzy memberships align well with the underlying cluster structure, reflecting compact and well-separated clusters with minimal ambiguity in membership assignments. Therefore, when no prior knowledge is available for a given dataset, we recommend performing a grid search. For example, $C\in\{2,3,\cdots,6\}$, as suggested in Ferraro et al. (2019), and $m\in[1.2,2.5]$ (a common range for the fuzziness parameter), and selecting the configuration that yields the highest FSI. However, as in our simulation and real data application, we focus on detecting two different brain states, i.e., drowsy and alert. Thus, we fix $C=2$ in our paper.
3 Simulation
@@
In this section, we assess the performance of FuzzCoh by generating heavy-tailed fuzzy multivariate time series based on the simulated examples in Talento et al. (2024). We begin with a mixture of second-order autoregressive processes (AR(2)) for different blocks, that is, for eight channels/time series ($p=4$ and $q=4$) we have
$\begin{pmatrix}\boldsymbol{X}^{(b)}(t)\\ \boldsymbol{Y}^{(b)}(t)\\ \end{pmatrix}=[\boldsymbol{A}_{1}D^{(b)}(t)+\boldsymbol{A}_{0}(1-D^{(b)}(t))]\boldsymbol{O}^{(b)}(t)+\boldsymbol{W}^{(b)}(t)$ (11)
@@
where $\boldsymbol{X}^{(b)}(t)\in\mathbb{R}^{p}$ and $\boldsymbol{Y}^{(b)}(t)\in\mathbb{R}^{q}$ are the concatenated multivariate time series at block $b$. The vector of latent processes, $\boldsymbol{O}^{(b)}(t)\in\mathbb{R}^{5}$, is composed of different mutually exclusive AR(2)’s. The parameters of ${O}_{r}^{(b)}(t)$, for $r=1,\dots,5$, are $\phi_{1,r}=2\exp(-M)cos(2\pi\omega_{r})\ \text{ and }\ \phi_{2}=-\exp(-2M)$, where $M=1.05$, $S\omega_{r}\in\{2,6,10,20,40\}$ and $S=128$ Hz (settings are followed from Talento et al., 2024). Observe that $S\omega_{r}$ belongs to any of the five frequency bands defined in Section 2 (details can be found in Ombao and Pinto, 2024).
@@
The fuzzy series are generated through binary indicator, $D^{(b)}(t)=c$, for $c\in\{0,1\}$. That is, we generate MTS with mixing-matrix $\boldsymbol{A}_{c}\in\mathbb{R}^{(p+q)\times 5}$ if time point, $t$, belongs to cluster-$c$. This formulation allows us to create blocks of multivariate time series that exclusively belong to one of the two clusters, as well as blocks that contain contributions from both processes, i.e., representing fuzzy series. The $b$-th multivariate time series is said to be pure if $D^{(b)}(1)=D^{(b)}(2)=\dots=D^{(b)}(T)=c$ and $c\in\{0,1\}$. The $b$-th multivariate time series is said to be fuzzy if $\mathbb{P}[D^{(b)}(t)=0]\in(0,1)$ and $\mathbb{P}[D^{(b)}(t)=1]=1-\mathbb{P}[D^{(b)}(t)=0]$ for all $t=1,\dots,T$.
@@
We provide here four examples that are different in terms of the independent white-noise processes, $\boldsymbol{W}(t)\in\mathbb{R}^{p+q}$. We let the marginal distribution of white noise processes to have the following distributions, namely, (i) Example 1: $W^{(b)}_{p}(t)\sim N(0,1)$; (ii) Example 2: $W^{(b)}_{p}(t)\sim\text{Student's }t_{3}$; and (iii) Example 3: $W^{(b)}_{p}(t)\sim\text{Student's }t_{1}$. Examples in (ii) and (iii) are known to be heavy-tailed and example in (iii) has non-existence of second moment. In this paper, a single block consists of 384 time points, i.e., 128 Hz sampling rate for 3 seconds. Moreover, we have $B=300$ yielding $115,200\times 8$ data points. The runtime for the whole framework in Figure 1 with this amount of data takes about 2 minutes. The codes are available in the Supplementary Material and implemented in parallel-run.
3.1 Comparison of methods and evaluation indicators
To show the performance of our proposed method, we include two additional methods for comparison.
@@ 3.1 Comparison of methods and evaluation indicators
    Fuzzy $C$-means (FCM) clustering. FCM is the extension of $K$‑means by allowing each observation to belong to every cluster with a graded membership (Bezdek et al., 1984). The method can be found in the R package [dtwclust], with the function [tsclust].
@@
    Variable-based Principal Component Analysis (VPCA) clustering. It reduces MTS data by applying PCA across the variable dimension, yielding a compact set of variable‑wise components that preserve spatial correlations for efficient downstream clustering. The method can be found in the R package [mlmts], with the function [vpca_clustering] (He and Tan, 2018).
@@
After assigning each block a clear label, we use the Rand index (RI) to compare the results of different methods. The RI measures the concordance between our clustering partitions and the true labels:
$\text{RI}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}},$ (12)
where TP and TN are the numbers of correctly clustered and correctly separated pairs, respectively, while FP and FN represent misclustered pairs. RI has range $[0,1]$. Higher RI values indicate greater similarity between the experimental and true partitions. This setup allows us to evaluate our proposed clustering method under both well-separated and partially overlapping time-series dynamics.
The results are shown in Figure 2, where we present the mean Rand Index (RI) of the three methods across 100 replications on the same dataset, using different fuzziness parameters. Since the FCM and VPCA methods cannot produce fuzzy clustering by band, we apply all three methods (FCM, FuzzCoh and VPCA) to the raw MTS without feature extraction from specific frequency bands.
Figure 2: The RI comparison
Figure 2 shows that FuzzCoh consistently maintained high accuracy, even in the presence of heavy-tailed noise, outperforming the other two methods. In contrast, FCM exhibited declining accuracy as noise distribution becomes more heavy, while VPCA performed the worst in terms of Rand Index (RI). Moreover, VPCA fails to converge in at least 50% of cases when the added noise follows a Cauchy distribution. These results highlight the robustness of FuzzCoh to extreme values and its effectiveness as a leading fuzzy clustering algorithm for multivariate time series.
4 Application
In this section, we evaluate the performance of our method using a real EEG dataset, which is publicly available at https://figshare.com/articles/dataset/EEG_driver_drowsiness_dataset/14273687?file=30707285. The dataset contains 2,022 trials of EEG recording collected from 11 individuals during a simulated driving task. Each trial consists of 384 time points (corresponding to 3 seconds at a sampling rate of 128 Hz), with each trial labeled as either drowsy or alert, i.e., $C=2$. We emphasize that our method accommodates more than two brain states; however, for the purposes of this application, we fix $C=2$. In our application, we defined four regions that are shown to be related with driving-task (Li et al., 2022; Talento et al., 2024), namely, (i) left frontal and pre-frontal lobe (LFp), (ii) right frontal and pre-frontal lobe (RFp), (iii) left temporo-parietal and occipital lobe (LTPO) and (iv) right temporo-parietal and occipital lobe (RTPO) (see leftmost panel of Figure 1). Our goal here is to address the following research questions:
@@ 4 Application
Table 1 presents the maximum RI per subject obtained under two settings: (i) using all connectivity vectors (i.e., all four brain regions predefined in Figure 1), within a specified frequency band, and (ii) selecting a pair of brain region within a specific frequency band that yields the highest accuracy per subject. The number of trials for each subject is also reported. The accuracy is evaluated against the labels defined by the experimenters based on the driver’s reaction time (see Cao et al., 2019, for details), which we treat as the ground truth. To have a label that can compare with the true label (i.e., ground truth), we assign the time series to the group using the maximum membership rule. For each configuration, we report the corresponding frequency band, accuracy, and the fuzziness parameter $m$ that yield the best results, together with the number of fuzziness time series we detected using a 0.7 threshold.
@@ 4.1 All Subject Analysis
The results that incorporate brain region connectivity provide additional interpretability by linking clustering metrics to underlying neural functions. Subject 6 and Subject 7 achieve their best accuracy using Theta band signals from the LFp–RTPO and LFp–RFp connections, respectively. This pattern highlights the pivotal role of the LFp in distinguishing between alert and drowsy states. The LFp is known to support working memory, executive functions, and language processing—core cognitive domains that are sensitive to fluctuations in mental alertness (Beauchamp, 2005). It plays a central role in goal-directed behavior, planning, and error monitoring, and is structurally connected to other regions such as the temporal and occipital cortices. These connections, particularly with RTPO (associated with sensory integration) and RFp (involved in cognitive control), may reflect the breakdown of coordinated top-down regulation as alertness declines. Thus, the prominence of LFp connectivity in achieving high clustering accuracy suggests that changes in executive and attentional networks are key markers of cognitive state transitions.
Table 1: Max RI per subject by frequency band and regional connectivity, with trial counts
@@
4.2 Individual Subject Analysis
We further investigate the signals for Subject 2 and Subject 6, which show relatively high RI, i.e., high separation of two states (see Table 1). Subject 2 has $m=1.2$ fuzziness parameter which indicates that the MTS are well-separable, while Subject 6 has $m=2.2$ fuzziness parameter for LFp-RTPO suggesting presence of fuzzy MTS.
Figure 3: The RI of all or partial brain regions, frequency band, and $m$ value of subject 2
Figure 4: Membership matrix of EEG samples collected from Subject 2 at Beta-band using LFp–LTPO connectivity
Figure 5: Discriminating functional brain connectivity structure (LFp–LTPO) of Subject 2 at Beta band when alert and when drowsy
Figure 6: (Top) Filtered signals and (bottom) summary of filtered signals from LTPO region during alert and drowsy states at Beta band of Subject 2
@@ 4.2 Individual Subject Analysis
Figure 7: The RI of all or partial brain regions, frequency band, and m value of Subject 6
Figure 8: Membership matrix of EEG samples collected from Subject 6 at Theta-band using LFp–RTPO connectivity
As suggested in Table 1, analyzing the regional connectivity, FuzzCoh suggests 8.43% of its EEG trails are fuzzy series. We would like to visualize both the discriminating and fuzzy scenarios. Figures 10 and 10 present the non-fuzzy observations and clearly highlight the distinctions between the two brain states. These observations accord with earlier work: Brown et al. (2012) reported a shift toward strong anterior Theta rhythms at the onset of drowsiness, and Sturm et al. (1999) identified evidence for a fronto-parietal-thalamic-brainstem network in the right hemisphere during alertness of 15 individuals
Figure 9: Discriminating functional brain connectivity structure (LFp–RTPO) of Subject 6 at Theta band when alert and when drowsy
Figure 10: (Top) Discriminating filtered signals from RTPO region during alert and drowsy states and (bottom) their corresponding summaries of Subject 6
Figure 11: Fuzzy functional brain connectivity structure of Subject 6 at Theta band when alert and when drowsy
Figure 12: (Top) Fuzzy filtered signals from RTPO region during alert and drowsy states and (bottom) their corresponding summaries of Subject 6
The most informative insight comes from the fuzzy cases. Figure 12 shows the fuzzy functional connectivity of Subject 6 in the Theta band. Unlike the discriminating trials, the two states display nearly identical connectivity: every node appears in both maps with comparable intensity, and no channel clearly dominates. Figure 12 plot the corresponding filtered time series for RTPO region. The waveforms and amplitudes are virtually indistinguishable between the alert (left) and drowsy (right) windows, underscoring the ambiguous nature of these trials. Their connectivity does not align decisively with either cluster, suggesting they represent a transitional or mixed brain state rather than a distinct condition.
@@ 5 Discussion
This paper introduced a robust spectral feature-based fuzzy clustering algorithm for multivariate time series, with a specific focus on EEG data. We began by introducing the concept of canonical variates and then formulated the clustering problem in detail. Through comparisons with two alternative methods across various (un)contamination scenarios, we demonstrated that our approach consistently outperforms the others in all cases. The main advantages of our FuzzCoh are: (1) resistant to EEG artifacts, (2) assigns graded memberships to capture the alert-to-drowsy transition, (3) operates in the frequency domain while explicitly modeling functional brain connectivity, and (4) by focusing on specific brain regions, our method performs dimension reduction, enabling clustering on high-dimensional time series data using only relevant channels. In short, by concentrating on a few discriminative frequency bands (Theta, Beta, and Gamma) and key fronto-parietal connections, our approach markedly sharpens the contrast between brain states, enabling earlier and more reliable drowsiness detection that strengthens driver-monitoring systems, enhances road safety, and benefits other spectral-spatial MTS applications.
Beyond its methodological strengths, our approach offers practical value across domains. In driver drowsiness detection, it enables early-warning systems by capturing subtle shifts in spectral and spatial EEG patterns (Stancin et al., 2021). In clinical diagnostics, it may help differentiate neurological disorders through changes in brain connectivity (Smith, 2005). For neuromarketing and cognitive workload assessment, it extracts interpretable features that reflect mental effort and engagement levels (Dehais et al., 2020).
Except for neuroscience, the FuzzCoh framework can be applied to financial time series by clustering coherence among asset groups to identify market regimes and systemic risk (Haldane and May, 2011; Biase and D’Apolito, 2012), or detect anomalies in cryptocurrency markets (Pocher et al., 2023) by capturing coordinated behavior across tokens at multiple frequencies. In environmental monitoring, it supports clustering of sensor networks (e.g., air quality or temperature) based on shared temporal patterns, enabling early detection of abnormal or synchronized environmental changes (Yin and Gaber, 2008).
By modeling both frequency- and region-specific dependencies within a robust fuzzy clustering framework, FuzzCoh serves as a versatile and interpretable tool for analyzing high-dimensional, multivariate systems in real-world contexts.
Despite its advantages, FuzzCoh remains sensitive to parameter tuning. It also assumes local stationarity, an assumption that may not hold in practice. Furthermore, its reliance on predefined regions and frequency bands can limit generalizability. Future work will therefore: (1) develop adaptive block segmentation and scalable approximations for large-scale MTS, (2) extend FuzzCoh to handle non-stationary MTS, and (3) generalize FuzzCoh for robust detection of outliers in MTS.

### v2

@@ 2.2.2 Cluster validity index for $C$ and $m$
A higher FSIΩ value indicates that the fuzzy memberships align well with the underlying cluster structure, reflecting compact and well-separated clusters with minimal ambiguity in membership assignments. Therefore, when no prior knowledge is available for a given dataset, we recommend performing a grid search. For example, $C\in\{2,3,\cdots,6\}$, as suggested in Ferraro et al. (2019); Ma et al. (2025), and $m\in[1.2,2.5]$ (a common range for the fuzziness parameter), and selecting the configuration that yields the highest FSI. However, as in our simulation and real data application, we focus on detecting two different brain states, i.e., drowsy and alert. Thus, we fix $C=2$ in our paper. Figure 1 shows the flow of the FuzzCoh algorithm.
Figure 1: The FuzzCoh framework.
3 Simulated EEG performance
This section evaluates the performance of FuzzCoh on both clean and contaminated MTS using the simulation designs of Talento et al. (2025). Our aim is to quantify accuracy on clean data and robustness under contamination, while benchmarking against representative alternatives.
3.1 Simulation design and settings
@@
We begin with a mixture of second-order autoregressive processes (AR(2)) for different blocks, that is, for eight channels/time series ($p=4$ and $q=4$) we have
$\bm{A}^{(b)}(t)=\bm{A}_{1}D^{(b)}(t)+\bm{A}_{0}(1-D^{(b)}(t)),$
$\begin{pmatrix}\bm{X}^{(b)}(t)\\ \bm{Y}^{(b)}(t)\\ \end{pmatrix}=\bm{A}^{(b)}(t)\bm{O}^{(b)}(t)+\bm{W}^{(b)}(t)$ (10)
@@
where $\bm{X}^{(b)}(t)\in\mathbb{R}^{p}$ and $\bm{Y}^{(b)}(t)\in\mathbb{R}^{q}$ are the concatenated MTS at block $b$. The vector of latent processes, $\bm{O}^{(b)}(t)\in\mathbb{R}^{5}$, is composed of different mutually exclusive AR(2)’s. The parameters of ${O}_{r}^{(b)}(t)$, for $r=1,\dots,5$, are $\phi_{1,r}=2\exp(-M)cos(2\pi\omega_{r})\ \text{ and }\ \phi_{2}=-\exp(-2M)$, where $M=1.05$, $S\omega_{r}\in\{2,6,10,20,40\}$ and $S=128$ Hz. Observe that $S\omega_{r}$ belongs to any of the five frequency bands defined in Section 2 (details can be found in Ombao and Pinto, 2024).
@@
The fuzzy series are generated through binary indicator, $D^{(b)}(t)=c$, for $c\in\{0,1\}$. That is, we generate MTS with mixing-matrix $\bm{A}_{c}\in\mathbb{R}^{(p+q)\times 5}$ if time point, $t$, belongs to cluster $c$. This formulation allows us to create blocks of MTS that exclusively belong to one of the two clusters, as well as blocks that contain contributions from both processes, i.e., representing fuzzy series. The $b$-th MTS is said to be pure if $D^{(b)}(1)=D^{(b)}(2)=\dots=D^{(b)}(T)=c$ and $c\in\{0,1\}$. The $b$-th MTS is said to be fuzzy if $\mathbb{P}[D^{(b)}(t)=0]\in(0,1)$ and $\mathbb{P}[D^{(b)}(t)=1]=1-\mathbb{P}[D^{(b)}(t)=0]$ for all $t=1,\dots,T$.
@@
We provide here four examples that are different in terms of the independent white-noise processes, $\bm{W}(t)\in\mathbb{R}^{p+q}$. We let the marginal distribution of white noise processes to have the following distributions, namely, (i) Example 1: $W^{(b)}_{p}(t)\sim N(0,1)$; (ii) Example 2: $W^{(b)}_{p}(t)\sim\text{Student's }t_{3}$; and (iii) Example 3: $W^{(b)}_{p}(t)\sim\text{Student's }t_{1}$. Examples in (ii) and (iii) are known to be heavy-tailed and example in (iii) has non-existence of second moment. In this paper, a single block consists of 384 time points, i.e., 128 Hz sampling rate for 3 seconds. Moreover, we have $B=300$ yielding $115,200\times 8$ data points. The runtime for the whole framework in Figure 1 with this amount of data takes about 2 minutes for a desktop with 3 gigahertz 6-Core Intel Core i5 processor and 16 gigabytes 2667 MHz DDR4 memory. The codes are available in the Supplementary Material and implemented in parallel-run.
3.2 Comparison methods and evaluation metrics
To evaluate the performance of the proposed method, we benchmark it against several alternatives. [^2] Code for the first three methods is in the R package mlmts; the FCPCA implementation is at https://github.com/arbitraryma/FCPCA.
@@ 3.2 Comparison methods and evaluation metrics
    Variable-based principal component analysis (VPCA) clustering. It reduces MTS data by applying PCA across the variable dimension, yielding a compact set of variable‑wise components that preserve spatial correlations for efficient downstream clustering (He and Tan, 2018).
@@
    Quantile cross-spectral density based fuzzy clustering (QCD). It captures complex dependence structures robustly, even under heavy-tailed distributions (López-Oriona et al., 2022).
  3. 3.
    Wavelet-based clustering of MTS (WAV). It decomposes each MTS into wavelet variances and wavelet correlations across multiple scales, capturing both variability within individual series and interactions between components (D’Urso and Maharaj, 2012).
  4. 4.
    Fuzzy clustering of MTS using common principal component analysis (FCPCA). estimating cross-covariance matrices within each cluster, extracting common principal components, and projecting series into low-dimensional subspaces for membership updates (Ma et al., 2025).
@@
After assigning each block a clear label, we use the Rand index (RI) to compare the results of different methods. The RI measures the concordance between our clustering partitions and the true labels. RI has range $[0,1]$. Higher RI values indicate greater similarity between the experimental and true partitions. This setup allows us to evaluate our proposed clustering method under both well-separated and partially overlapping time-series dynamics.
The results are shown in Figure 2, 3, and 4, where we present the mean Rand Index (RI) of all methods across 100 replications on the same dataset, using different fuzziness parameters. To have a fair comparison, we apply all methods to the raw MTS without feature extraction from specific frequency bands. Across all scenarios, FuzzCoh consistently outperforms the competing methods. It maintains high accuracy under Normal and Student’s t settings and is uniquely robust under the most contaminated case (Cauchy), where all other methods deteriorate severely. This demonstrates that FuzzCoh is stable across different noise regimes and resilient to heavy-tailed contamination.
In terms of efficiency shown in Figure 5, FuzzCoh runs within a few seconds, comparable to the fastest baselines and more than 20 times faster than QCD. Taken together, these results show that FuzzCoh achieves the best trade-off between accuracy, robustness, and runtime efficiency, making it the most reliable method across diverse conditions.
Figure 2: Simulation result under Example 1.
Figure 3: Simulation result under Example 2.
Figure 4: Simulation result under Example 3.
Figure 5: Average times comparisons of all methods under each example.
4 Real EEG data application
In this section, we evaluate the performance of our method using a real EEG dataset [^3] This dataset is available at https://figshare.com/articles/dataset/EEG_driver_drowsiness_dataset/14273687?file=30707285. The dataset contains 2,022 trials of EEG recording collected from 11 individuals during a simulated driving task. Each trial consists of 384 time points (corresponding to 3 seconds at a sampling rate of 128 Hz), with each trial labeled as either drowsy or alert, i.e., C=2C=2.444The RI for other methods are shwon in Appendix, Section 6.2. We emphasize that our method accommodates more than two brain states; however, for the purposes of this application, we fix $C=2$. In our application, we defined four regions that are shown to be related with driving-task (Li et al., 2022; Talento et al., 2025), namely, (i) left frontal and pre-frontal lobe (LFp), (ii) right frontal and pre-frontal lobe (RFp), (iii) left temporo-parietal and occipital lobe (LTPO) and (iv) right temporo-parietal and occipital lobe (RTPO) (see leftmost panel of Figure 1). Our goal here is to address the following research questions:
@@ 4 Real EEG data application
Table 1 presents the maximum RI per subject obtained under two settings: (i) using all connectivity vectors (i.e., all four brain regions predefined in Figure 1), within a specified frequency band, and (ii) selecting a pair of brain region within a specific frequency band that yields the highest accuracy per subject. The number of trials for each subject is also reported. The accuracy is evaluated against the labels defined by the experimenters based on the driver’s reaction time (see Cao et al., 2019, for details), which we treat as the ground truth. To have a label that can compare with the true label (i.e., ground truth), we assign the time series to the group using the maximum membership rule. For each configuration, we report the corresponding frequency band, accuracy, and the fuzziness parameter $m$ that yield the best results, together with the number of fuzziness time series we detected using a 0.7 threshold (see (D’Urso and Maharaj, 2009; López-Oriona et al., 2022; Ma et al., 2025) for a detailed discussion of the threshold).
@@ 4.1 All subject analysis
The results that incorporate brain region connectivity provide additional interpretability by linking clustering metrics to underlying neural functions. Subj6 and Subj7 achieve their best accuracy using Theta band signals from the LFp–RTPO and LFp–RFp connections, respectively. This pattern highlights the pivotal role of the LFp in distinguishing between alert and drowsy states. The LFp is known to support working memory, executive functions, and language processing—core cognitive domains that are sensitive to fluctuations in mental alertness (Beauchamp, 2005). It plays a central role in goal-directed behavior, planning, and error monitoring, and is structurally connected to other regions such as the temporal and occipital cortices. These connections, particularly with RTPO (associated with sensory integration) and RFp (involved in cognitive control), may reflect the breakdown of coordinated top-down regulation as alertness declines. Thus, the prominence of LFp connectivity in achieving high clustering accuracy suggests that changes in executive and attentional networks are key markers of cognitive state transitions.
Table 1: Max RI per subject by frequency band and regional connectivity, with trial counts.
@@
4.2 Selected individual subject analysis
We further investigate the signals for Subject 2 and Subject 6, which show relatively high RI, i.e., high separation of two states (see Table 1). Subject 2 has $m=1.2$ fuzziness parameter which indicates that the MTS are well-separable, while Subject 6 has $m=2.2$ fuzziness parameter for LFp-RTPO suggesting presence of fuzzy MTS  [^5] We show in Appendix, Section 6.3 how to use FSI to guide the selection of hyperparamters in case of no prior infomation of the dataset is known. .
Figure 6: The RI of all or partial brain regions, frequency band, and $m$ value of subject 2.
Figure 7: Membership matrix of EEG samples collected from Subject 2 at Beta-band using LFp–LTPO connectivity.
Figure 8: Functional brain connectivity structure (LFp–LTPO) of Subject 2 at Beta band when alert and when drowsy.
Figure 9: Filtered signals from LTPO region during alert and drowsy states at Beta band of Subject 2.
@@ 4.2 Selected individual subject analysis
Figure 10: The RI of all or partial brain regions, frequency band, and m value of Subject 6.
Figure 11: Membership matrix of EEG samples collected from Subject 6 at Theta-band using LFp–RTPO connectivity.
As suggested in Table 1, analyzing the regional connectivity, FuzzCoh suggests 8.43% of its EEG trails are fuzzy series. We would like to visualize both the discriminating and fuzzy scenarios. Figures 12 (top) and 13 present the non-fuzzy observations and clearly highlight the distinctions between the two brain states. These observations accord with earlier work: Brown et al. (2012) reported a shift toward strong anterior Theta rhythms at the onset of drowsiness, and Sturm et al. (1999) identified evidence for a fronto-parietal-thalamic-brainstem network in the right hemisphere during alertness of 15 individuals.
The most informative insight comes from the fuzzy cases. Figure 12 (bottom) shows the fuzzy functional connectivity of Subject 6 in the Theta band. Unlike the discriminating trials, the two states display nearly identical connectivity: every node appears in both maps with comparable intensity, and no channel clearly dominates.
Figure 13 plot the corresponding filtered time series for RTPO region. The fuzzy waveforms and amplitudes are virtually indistinguishable between the alert (left) and drowsy (right) windows, underscoring the ambiguous nature of these trials. Their connectivity does not align decisively with either cluster, suggesting they represent a transitional or mixed brain state rather than a distinct condition.
Figure 12: (Top) Discriminating and (bottom) fuzzy functional brain connectivity structure (LFp–RTPO) of Subject 6 at Theta band when alert and when drowsy.
Figure 13: (Top) Discriminating and (bottom) fuzzy filtered signals from RTPO region during alert and drowsy states of Subject 6.
@@ 5 Discussion
This paper introduces a robust spectral feature-based fuzzy clustering algorithm for MTS, with EEG data serving as a motivating application. We first outline the concept of canonical variates and then formulate the clustering problem in detail. Through extensive comparisons with alternative methods under both clean and contaminated conditions, we show that our approach consistently achieves superior performance. The main advantages of the proposed FuzzCoh method are: (1) robustness to outliers, (2) graded memberships that capture overlapping cluster structure, (3) operation in the frequency domain with explicit modeling of functional connectivity across multiple locations or variables, and (4) dimension reduction via region-specific connectivity patterns, enabling effective clustering of high-dimensional MTS data using only the most informative features. Beyond EEG, the method is broadly applicable to other spectral–spatial MTS domains.
In terms of practical impact, our approach provides value across diverse fields. In driver drowsiness detection, it enables early-warning systems by identifying subtle changes in spectral and spatial EEG patterns (Stancin et al., 2021). In clinical diagnostics, it can aid the differentiation of neurological disorders through connectivity alterations (Smith, 2005). For neuromarketing and cognitive workload assessment, it extracts interpretable features reflecting mental effort and engagement (Dehais et al., 2020).
Beyond neuroscience, the FuzzCoh framework extends naturally to other applications. In finance, clustering coherence among asset groups can reveal market regimes and systemic risk (Haldane and May, 2011; Biase and D’Apolito, 2012), while coherence patterns across cryptocurrencies can be used to detect anomalies and coordinated behavior at multiple frequencies (Pocher et al., 2023). In environmental monitoring, it facilitates the clustering of sensor networks (e.g., air quality or temperature) based on shared temporal patterns, enabling early detection of abnormal or synchronized environmental changes (Yin and Gaber, 2008).
By modeling both frequency- and region-specific dependencies within a robust fuzzy clustering framework, FuzzCoh serves as a versatile and interpretable tool for analyzing high-dimensional MTS in real-world contexts.
Despite these strengths, FuzzCoh has limitations. Its performance depends on careful parameter tuning, and it assumes local stationarity, an assumption that may not always hold. Moreover, reliance on predefined regions and frequency bands may restrict its adaptability in certain applications. Future work will therefore focus on: (1) developing adaptive block segmentation and scalable approximations for large-scale MTS, (2) extending FuzzCoh to non-stationary settings, and (3) generalizing the framework for robust outlier detection in MTS.